OpenStudy (anonymous):

Find The GCF of 18x3 and 30x5

OpenStudy (anonymous):

GCF can be find by finding common terms of the abov terms x^3 *6 18x3=2*3*3*x^3 30x5=2*3*5*x^3*x^2 the common terms are 2*3*x^2=6x2

OpenStudy (anonymous):

Not quite @sai1234 .

OpenStudy (anonymous):

54 150 2* 27 75 3* 9 25 3 3 25 3 1 25 25 * : common divides so 2*3 = 6 = GCF

OpenStudy (anonymous):

Both of you have the numeral part of the GCF correct, but not the variable portion. \(x^3 = x \times x \times x\) \(x^5 = x \times x \times x \times x \times x\)You can factor out an x^3 since it is in both. THerefore, it is: \(GCF = 6x^3\)